One mechanic services 8 drilling machines for a steel plate manufacturer. Machines break down on an average of once every 8 working​ days, and breakdowns tend to follow a Poisson distribution. The mechanic can handle an average of one repair job per day. Repairs follow a negative exponential distribution.

​a) On the​ average, how many machines are waiting for​ service?

The average number of machines waiting for service is _. (Round your response to three decimal​ places.)

​b) On the​ average, what is the waiting time to be​ serviced?

The average waiting time to be serviced is _ days. ​(Round your response to two decimal​ places.)
Answer #1

a) Machines waiting for service

Avg of machine breakdowns in a day = 1/8 = 0.125

Avg of repairs = 1

L = 0.125 /(1-0.125) = 0.143

Thus, avg number of machines waiting = 0.143

b) Waiting time , W

= 0.125/1 (0.875) = 0.14 days